# Detecting the tidal heating with the generic extreme-mass-ratio inspirals

### v2

@@ II.4 Waveform computation and analysis
The few model is a fast waveform generation tools for the future EMRI data analysis, which includes the eccentric adiabatic orbits for the Schwarzschild BH and the generic adiabatic orbits at 5PN order for the Kerr BH Katz et al. (2021, 2023); Chua et al. (2017); Speri et al. (2023). We compute the waveforms in the few model using the inspiral trajectories of EMRI modified by the tidal heating effect as introduced in the subsection II.3. With the EMRI trajectories produced by the few model, we can assess the effect of tidal heating on waveform phase by computing the dephasing of EMIR waveforms. The waveform dephasing can be defined by the integral of the differences of orbital frequencies between the Kerr BH case and the ECO case
@@
where $\Omega_{i}\in(\Omega_{r},\Omega_{\theta},\Omega_{\phi})$ are the orbital fundamental frequencies that are derived by Refs. Schmidt (2002); Fujita and Hikida (2009). The superscripts $\mathcal{R}=0$ and $\mathcal{R}\neq 0$ correspond to the frequencies computed with the trajectories of the Kerr BH and the ECO, respectively. A useful criterion for the detection of additional effects on waveforms has been proposed in Refs.Datta et al. (2020); Speeney et al. (2022), where the effect from tidal heating can be detected if the mismatch satisfies $\delta\Phi_{i}\geq 1$.
To quantitatively assess the effect of tidal heating on EMRI waveforms with space-borne GW detectors, we can perform the mismatch analysis using two kinds of waveform from EMRI system with and without the tidal heating effect. For two waveforms $h_{a}(t)$ and $h_{b}(t)$, the mismatch $\mathcal{M}$ can be defined by
$\mathcal{O}=1-\frac{<h_{a}|h_{b}>}{\sqrt{<h_{a}|h_{a}><h_{b}|h_{b}>}},$ (24)
and the noise-weighted inner product <*|*><*|*>< * | * > is defined by
@@
where the quantities with tildes denote the Fourier transform of waveform data, the star signs mean complex conjugation, and $S_{n}(f)$ is noise power spectral density of LISA Amaro-Seoane et al. (2017). In this paper, we adopt a useful formula to distinguish between two waveforms in LISA observations. LISA can detect the difference in waveforms if the mismatch satisfies $\mathcal{M}\geq 1/(2\rho^{2})$ Flanagan and Hughes (1998); Lindblom et al. (2008). The minimum signal-to-noise ratio (SNR) of the EMRI signal detected by LISA or TianQin is $\rho=20$, so the threshold value of mismatch is $\mathcal{M}_{\rm min}=0.001$.
To assess the capability of future space-based GW detectors, such as LISA, for measuring the parameter $|\mathcal{R}|^{2}$ of tidal heating, we plan to perform the parameter estimation using statistic methods. The FIM is a common method to estimate the measure error of GW source parameters, which can be defined by
$\Gamma_{ij}=\left(\frac{\partial h_{i}}{\partial\theta_{i}}|\frac{\partial h_{j}}{\partial\theta_{j}}\right)$ (26)
where $\theta_{i}$, $i=1,2,3,\cdots$, are the parameters of the modified few model describing EMRI systems with ECO. The intrinsic parameters are classified as $\theta_{\rm intr}=(M,m,a,p,e,x_{I},|\mathcal{R}|^{2},\Phi_{\phi}^{0},\Phi_{\theta}^{0},\Phi_{r}^{0})$, where the quanties $(\Phi_{\phi}^{0},\Phi_{\theta}^{0},\Phi_{r}^{0})$ represent for the initial values of azimuthal, polar and radial phases, and the extrinsic parameters are $\theta_{\rm extr}=(D,\vartheta_{S},\varphi_{S},\vartheta_{K},\varphi_{K})$. $D$ refers to the distance from the source to the detector, and $(\vartheta_{S},\varphi_{S},\vartheta_{K},\varphi_{K})$ are the azimuth and colatitude of source sky position and orbital angular momentum. For the EMRI GW signal with higher SNR, the uncertainty of $i-\rm th$ parameter $\theta_{i}$ can be obtained by the square root of the variance-covariance matrix
$\sigma_{i}=\Delta\theta_{i}=\Sigma_{aa}^{1/2}$ (27)
where the variance-covariance matrix is obtained by the inverse of the FIM
$\Sigma_{ab}=(\Gamma_{ab})^{-1}.$ (28)
It should be noted that the precondition of FIM method is linear signal approximation, whose applicability has been argued in Ref. Zi et al. (2023).
@@
$\Delta\Omega_{i}=2\pi|\sin\vartheta_{i}|\sqrt{\sigma^{2}_{\vartheta_{i}}\sigma^{2}_{\varphi_{i}}-\Sigma^{2}_{\vartheta_{i}\varphi_{i}}}$ (29)
with $i=(S,K)$.
@@ III.1 Spinning massive objects
Firstly, we present a comparison of EMRI waveforms generated with the modified few model. In this plot, the massive object has a mass of $10^{6}M_{\odot}$, a spin of $a=0.9$, a reflectivity of $|\mathcal{R}|^{2}=0.9$, the smaller object has a mass of $30M_{\odot}$, and the distance of the EMRI is $D_{L}=1\rm Gpc$. Note that the subsequent figures retain the same configuration. As illustrated in Fig. 1, initially, the phases of the two EMRI waveforms closely coincide in the left panel of the top figure. However, as time progresses, specifically two months later, distinct disparities in the phases of the two EMRI waveforms become more pronounced in the right panel of the top figure. The bottom of the figure displays the complete time domain waveform within one year. Based on Figure 1, it can be inferred that the influence of tidal heating gradually intensifies over time.
@@
Figure 1: Comparison of the time domain EMRI waveforms from the Kerr BH and ECO with a reflectivity $|\mathcal{R}|^{2}=0.9$ and spin $a=0.9$, where the initial orbital parameters are $p_{0}=12,e_{0}=0.5,x_{I}=0.5$.
@@
Figure 2: The azimuthal dephasing $\delta\Phi_{\phi}$ (top left panel), radial dephasing $\delta\Phi_{r}$ (top right panel) and polar dephasing $\delta\Phi_{\theta}$ of the EMRI waveforms as a function of observation time with various values of the reflectivity $|\mathcal{R}|^{2}\in(10^{-5},5\times 10^{-5},10^{-4},10^{-3},10^{-2},10^{-1},0.3)$ are depicted, where the initial orbital parameters are $p_{0}=12,e_{0}=0.5,x_{I}=0.5$.
@@
The azimuthal, radial and polar dephasing of EMRI waveforms, as a function of observation time, is plotted in Fig. 2, where the initial orbital parameters are set the same as the ones in Fig. 1. The black and dashed horizontal lines in Fig. 2 represent the threshold values $\delta\Phi_{i}\sim 1\rm rad$. This is a rough constraint on the tidal heating effect, and the validity of this standard has been argued in Ref. Datta et al. (2020). All the three panels for the dephasings of different coordinates are confirmed to be capable of exceeding the threshold value when the suitable parameters are taken. One can see that the accumulated dephasing varies from the fraction radians to $10^{2}$ radians as a whole, particularly for the azimuthal dephasing $\Omega_{\phi}$. As argued in the Ref. Barsanti et al. (2022); Gupta et al. (2021), the azimuthal dephasing plays the dominant role compared to the dephasings of the other coordinates. According to Fig. 2, the radial and polar dephasings are both less than the azimuthal dephasing, which is consistent with the results in the Ref. Barsanti et al. (2022); Gupta et al. (2021). After considering the proper parameters, the accumulative effect of radial dephasing could be discernible. In summary, the tidal heating effect can lead to a significant detectability in the EMRI observations.
@@
Though the aforementioned dephasing results could help to understand the detectable effect of tidal heating on EMRI waveforms, we cannot claim that the minimum value of reflectivity can be distinguished with the LISA observation. Thus, we perform the mismatch analysis to quantitatively assess the corrected effects of tidal heating on EMRI waveforms. In Fig. 3, the mismatches as a function of observation time for different values of the reflectivity and the initial orbital eccentricity are plotted. Note that the horizontal black dash line denotes to the threshold value of mismatch can be discerned by LISA. As a whole, after setting proper initial conditions, the mismatch can be larger than the threshold value, which implies that the tidal heating effect is detectable in the EMRI observations. According to mismatches results in Ref. Datta et al. (2020), EMRI waveforms from the circular and equatorial orbits near the ECO with a reflectivity parameter $|\mathcal{R}|^{2}=10^{-4}$ can be just distinguished with thirteen months observation of LISA. From the left panel of Fig. 3, however, the EMRI waveform corrected by tidal heating effect, emitting from the eccentrical EMRI orbits, can be recognized for the initial orbital eccentric $e_{0}=0.5$ with about seven months observation of LISA. As shown in the left panel, the mismatch increases with the reflectivity parameter $|\mathcal{R}|^{2}$. For the right panel, the similar conclusions can be obtained if the initial orbital eccentricity is increasing. For a given initial eccentricity $e_{0}=0.4$ and $|\mathcal{R}|^{2}=5\times 10^{-4}$, LISA can discern the tidal heating effect through a five-month observation period.
@@
Figure 3: Mismatch as a function of observation time for EMRI waveforms from the ECO and the Kerr BH are depicted. The left panel shows the effects from various reflectivity parameters $|\mathcal{R}|^{2}\in(10^{-6},5\times 10^{-5},10^{-4},5\times 10^{-4},10^{-3},5\times 10^{-3},10^{-2},0.05,0.1)$ with $e_{0}=0.5$, and the right panel focus on the effects of orbital eccentricities $e_{0}\in(0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9)$ with $|\mathcal{R}|^{2}=5\times 10^{-4}$ on the mismatch. The other parameters are set as $a=0.9,p_{0}=12,x_{I}=0.5$.
@@
$p_{0}$ $e_{0}$ $\Delta M/M$ $\Delta m/m$ $\Delta a$ $\Delta e_{0}$ $\Delta p_{0}$ $\Delta x_{I}$ $\Delta(|\mathcal{R}|^{2})$ $\Delta D_{L}/D_{L}$ $\Delta\Omega_{S}$ $\Delta\Omega_{K}$
$12$ $0.4$ $3.44\times 10^{-5}$ $9.88\times 10^{-5}$ $1.71\times 10^{-4}$ $7.68\times 10^{-5}$ $8.71\times 10^{-4}$ $8.71\times 10^{-4}$ $5.07\times 10^{-4}$ $7.58\times 10^{-2}$ $6.09\times 10^{-4}$ $5.62$
$12$ $0.8$ $2.93\times 10^{-5}$ $6.04\times 10^{-5}$ $8.28\times 10^{-5}$ $5.21\times 10^{-5}$ $6.26\times 10^{-4}$ $8.43\times 10^{-4}$ $2.09\times 10^{-4}$ $4.69\times 10^{-2}$ $5.82\times 10^{-4}$ $4.35$
Table 1: Measurement errors of the EMRI source parameters including the reflectivity are listed, the other parameters are set as $\Phi_{r}^{0}=1.0$, $\Phi_{\phi}^{0}=1.0$ and $|\mathcal{R}|^{2}=0.001$.
Finally, we calculate the accuracy of source parameters, including reflectivity, using the FIM method. The EMRI waveforms are generated by the few model. In Table 1, we list the measurement errors on the source parameters with the LISA observations when the tidal heating effect is considered. From the table, we find that the reflectivity parameter $|\mathcal{R}|^{2}$ of the ECO can be constrained within a fractional error of ${}10^{-4}$ via the observations of LISA. The measurement errors of the full waveform parameters is slightly decreased with the initial orbital eccentricity.

### v6

@@ II.4 Waveform computation and analysis
The FEW model is a fast waveform generation tools for the future EMRI data analysis, which includes the eccentric adiabatic orbits for the Schwarzschild BH and the generic adiabatic orbits at 5PN order for the Kerr BH Katz:2021yft ; Katz et al. (2023); Chua et al. (2017); Speri et al. (2023). We compute the waveforms in the FEW model using the inspiral trajectories of EMRI modified by the tidal heating effect as introduced in the subsection II.3. With the EMRI trajectories produced by the FEW model, we can assess the effect of tidal heating on waveform phase by computing the dephasing of EMIR waveforms. The waveform dephasing can be defined by the integral of the differences of orbital frequencies between the Kerr BH case and the ECO case
@@
where $\Omega_{i}\in\{\Omega_{r},\Omega_{\theta},\Omega_{\phi}\}$ are the orbital fundamental frequencies that are derived by Refs. Schmidt (2002); Fujita and Hikida (2009). The superscripts $\mathcal{R}=0$ and $\mathcal{R}\neq 0$ correspond to the frequencies computed with the trajectories of the Kerr BH and the ECO, respectively. A useful criterion for the distinguishing additional effects on waveforms has been proposed in Refs.Datta et al. (2020); Speeney et al. (2022), where the effect from tidal heating can be discerned if the dephasing satisfies $\delta\Phi_{i}\geq 1$. Note that as pointed out by Ref. Datta et al. (2020) the dephasing only provides a very rough evaluation for the excluding parameters of tidal heating and a more careful analysis is needed.
To better assess the effect of tidal heating on EMRI waveforms with space-borne GW detectors, we perform the mismatch analysis using two kinds of waveform where one is the Kerr waveform and the other is the ECO’s. For two waveforms $h_{a}(t)$ and $h_{b}(t)$, the mismatch $\mathcal{M}$ is defined by
$\mathcal{M}=1-\frac{<h_{a}|h_{b}>}{\sqrt{<h_{a}|h_{a}><h_{b}|h_{b}>}},$ (24)
and the noise-weighted inner product <|><*|*>< ∗ | ∗ > is defined by
@@
where the quantities with tildes denote the Fourier transform of waveform data, the star signs mean complex conjugation, and $S_{n}(f)$ is noise power spectral density of LISA Amaro-Seoane et al. (2017). The signal-to-noise ratio (SNR) can be defined by
$\rho=\sqrt{<h|h>}.$ (26)
From the definition of mismatch in Eq. (24), one can see that the mismatch is zero for two identical waveforms. In this paper, we adopt a useful rule of thumb to distinguish between two waveforms in LISA observations. LISA can distinguish the differences in waveforms if the mismatch satisfies $\mathcal{M}\geq 1/(2\rho^{2})$ Flanagan and Hughes (1998); Lindblom et al. (2008). The SNR threshold for EMRI signals detected by the LISA or TianQin is chosen as $\rho=20$ Babak et al. (2017); Fan et al. (2020), so the threshold value of mismatch is $\mathcal{M}_{\rm min}=0.001$. Note that when calculating the mismatch we only vary the reflectivity and let all other parameters fixed. Since the those other parameters will be unknown a priori, this may lead to a higher mismatch than the realistic case.
To assess the capability of future space-based GW detectors, such as LISA, for measuring the parameter $|\mathcal{R}|^{2}$ of tidal heating, we perform the parameter estimation using statistic methods in this paper. The FIM is a common method to estimate the measure error of GW source parameters, which can be defined by
$\Gamma_{ij}=\left(\frac{\partial h}{\partial\theta_{i}}|\frac{\partial h}{\partial\theta_{j}}\right)$ (27)
where $\theta_{i}$, $i=1,2,3,\cdots$, are the parameters of the modified few model describing EMRI systems with ECO. The intrinsic parameters are classified as $\theta_{\rm intr}=(M,m,a,p_{0},e_{0},x_{I},|\mathcal{R}|^{2},\Phi_{\phi}^{0},\Phi_{\theta}^{0},\Phi_{r}^{0})$, where the quanties $(\Phi_{\phi}^{0},\Phi_{\theta}^{0},\Phi_{r}^{0})$ represent for the initial values of azimuthal, polar and radial phases, and the extrinsic parameters are $\theta_{\rm extr}=(D_{L},\vartheta_{S},\varphi_{S},\vartheta_{K},\varphi_{K})$. $D_{L}$ refers to the luminosity distance from the source to the detector, and $(\vartheta_{S},\varphi_{S},\vartheta_{K},\varphi_{K})$ are the azimuth and colatitude of source sky position and orbital angular momentum. For the EMRI signal with higher SNR, the variance-covariance matrix $\mathbf{\Sigma}$ can be given by the inverse of FIM
$<\delta\theta^{i}\delta\theta^{j}>\backsimeq(\mathbf{\Gamma}^{-1})^{ij}=\mathbf{\Sigma}^{ij},$ (28)
the uncertainty of $i-\rm th$ parameter $\theta_{i}$ can be obtained by the square root of the variance-covariance matrix
$\Delta\theta_{i}=\sqrt{<(\delta\theta_{i})^{2}>}\simeq\sqrt{(\mathbf{\Sigma}_{ii})}.$ (29)
It should be noted that the precondition of FIM method is linear signal approximation, whose applicability has been argued in Ref. Zi et al. (2023). In principle, the FIM method is more rigorous and reliable than the previous two method to assess the impact of tidal heating on the waveform. For example, the correlation in parameters can be described by the off-diagonal elements of the inverse FIM. Moreover, the FIM can embody the impact of the reflectivity on the measurement accuracy of the other parameters.
@@
$\Delta\Omega_{i}=2\pi|\sin\vartheta_{i}|\sqrt{\mathbf{\Sigma}^{2}_{\vartheta_{i}}\mathbf{\Sigma}^{2}_{\varphi_{i}}-\mathbf{\Sigma}^{2}_{\vartheta_{i}\varphi_{i}}}$ (30)
with $i\in\{S,K\}$.
@@ III.1 Spinning massive objects
Firstly, we present a comparison of EMRI waveforms generated with the modified FEW model. In all the following plots, the massive object has a mass of $10^{6}M_{\odot}$, a spin of $a=0.9$, a reflectivity of $|\mathcal{R}|^{2}=0.9$, the smaller object has a mass of $30M_{\odot}$, and the distance of the EMRI is $D_{L}=1\rm Gpc$. Note that the subsequent figures retain the same configuration. As illustrated in Fig. 1, initially, the phases of the two EMRI waveforms closely coincide in the left panel of the top figure. However, as time progresses, specifically two months later, distinct disparities in the phases of the two EMRI waveforms become more pronounced in the right panel of the top figure. The bottom of the figure displays the complete time domain waveform within one year. Based on the left panel and right panel of upper part in the Fig. 1, one can find that the GW phases for the Kerr cases keep same with the phases for the EMRI system with the tidal heating parameters $|\mathcal{R}|^{2}\in\{0.1,0.3,0.5,0.9\}$, after about two months the phases differences between the Kerr case and the ECO’ s case gradually increase over time.
@@
Figure 1: Comparison of the time domain EMRI waveforms from the Kerr BH and ECO with a reflectivity $|\mathcal{R}|^{2}=0.9$ and spin $a=0.9$, where the initial orbital parameters are $p_{0}=12,e_{0}=0.5,\theta_{I}=\pi/3$.
@@
Figure 2: The azimuthal dephasing $\delta\Phi_{\phi}$ (top left panel), radial dephasing $\delta\Phi_{r}$ (top right panel) and polar dephasing $\delta\Phi_{\theta}$ of the EMRI waveforms as a function of observation time with various values of the reflectivity $|\mathcal{R}|^{2}\in\{10^{-5},5\times 10^{-5},10^{-4},10^{-3},10^{-2},10^{-1},0.3\}$ are depicted, where the initial orbital parameters are $p_{0}=12,e_{0}=0.5,\theta_{I}=\pi/3$.
@@
The azimuthal, radial and polar dephasing of EMRI waveforms, as a function of observation time, is plotted in Fig. 2, where the initial orbital parameters are set the same as the ones in Fig. 1. The black and dashed horizontal lines in Fig. 2 represent the threshold values $\delta\Phi_{i}\sim 1\rm rad$. This provides a very rough constraint on the tidal heating effect, and the shortcoming of this standard has been argued in Ref. Datta et al. (2020). All the three panels for the dephasings of different coordinates are confirmed to be capable of exceeding the threshold value when the suitable parameters are taken. One can see that the accumulated dephasing varies from the fraction radians to $10^{2}$ radians as a whole, particularly for the azimuthal dephasing $\Omega_{\phi}$. As argued in the Ref. Barsanti et al. (2022); Gupta et al. (2021), the azimuthal dephasing plays the dominant role compared to the dephasings of the other coordinates. According to Fig. 2, the radial and polar dephasings are both less than the azimuthal dephasing, which is consistent with the results in the Ref. Barsanti et al. (2022); Gupta et al. (2021). After considering the proper parameters, the accumulative effect of radial dephasing could be discernible. In summary, the tidal heating effect can lead to a significant distinguishing criterion in the EMRI observations.
@@
Though the aforementioned dephasing results could give the standard that identified the effect of tidal heating on EMRI waveforms, we cannot claim that the minimum value of reflectivity can be distinguished with the LISA observation. Thus, we perform the mismatch analysis using two kinds of waveforms from the EMRI systems where one is the Kerr waveform and the other is the ECO’s. In Fig. 3, the mismatches as a function of observation time for different values of the reflectivity and the initial orbital eccentricity are plotted. Note that the horizontal black dashed line denotes to the threshold value of mismatch can be discerned by LISA. As a whole, after setting proper initial conditions, the mismatch can be larger than the threshold value, which implies that the tidal heating effect is distinguished in the EMRI observations. According to mismatches results in Ref. Datta et al. (2020), EMRI waveforms from the circular and equatorial orbits near the ECO with a reflectivity parameter $|\mathcal{R}|^{2}=10^{-4}$ can be just distinguished with thirteen months observation of LISA. From the left panel of Fig. 3, however, the EMRI waveform corrected by tidal heating effect, emitting from the eccentrical EMRI orbits, can be recognized for the initial orbital eccentric $e_{0}=0.5$ with about seven months observation of LISA. As shown in the left panel, the mismatch increases with the reflectivity parameter $|\mathcal{R}|^{2}$. For the right panel, the similar conclusions can be obtained if the initial orbital eccentricity is increasing. For a given initial eccentricity $e_{0}=0.4$ and $|\mathcal{R}|^{2}=5\times 10^{-4}$, LISA can discern the tidal heating effect through a five-month observation period.
@@
Figure 3: Mismatch as a function of observation time for EMRI waveforms from the ECO and the Kerr BH are depicted. The left panel shows the effects from various reflectivity parameters $|\mathcal{R}|^{2}\in\{10^{-6},5\times 10^{-5},10^{-4},5\times 10^{-4},10^{-3},5\times 10^{-3},10^{-2},0.05,0.1\}$ with $e_{0}=0.5$, and the right panel focus on the effects of orbital eccentricities $e_{0}\in\{0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9\}$ with $|\mathcal{R}|^{2}=5\times 10^{-4}$ on the mismatch. The other parameters are set as $a=0.9,p_{0}=12,\theta_{I}=\pi/3$.
@@
$p_{0}$ $e_{0}$ $\Delta\ln M$ $\Delta\ln m$ $\Delta a$ $\Delta e_{0}$ $\Delta p_{0}$ $\Delta x_{I}$ $\Delta(|\mathcal{R}|^{2})$ $\Delta D_{L}/D_{L}$ $\Delta\Omega_{S}$ $\Delta\Omega_{K}$
$12$ $0.4$ $2.35\times 10^{-5}$ $1.23\times 10^{-5}$ $1.75\times 10^{-6}$ $1.39\times 10^{-6}$ $2.18\times 10^{-5}$ $2.78\times 10^{-6}$ $1.82\times 10^{-6}$ $1.82\times 10^{-2}$ $6.09\times 10^{-4}$ $4.45$
$12$ $0.8$ $1.43\times 10^{-5}$ $6.04\times 10^{-7}$ $8.28\times 10^{-7}$ $1.21\times 10^{-6}$ $6.26\times 10^{-4}$ $1.43\times 10^{-6}$ $2.09\times 10^{-6}$ $4.69\times 10^{-2}$ $5.82\times 10^{-4}$ $2.75$
Table 1: Measurement errors of the EMRI source parameters including the reflectivity are listed, the other parameters are set as $\Phi_{r}^{0}=1.0$, $\Phi_{\theta}^{0}=1.0$, $\Phi_{\phi}^{0}=1.0$ and $|\mathcal{R}|^{2}=0.001$.
Finally, we calculate the accuracy of source parameters, including reflectivity, using the FIM method. The EMRI waveforms are generated by the FEW model. In Table 1, we list the measurement errors on the source parameters with the LISA observations when the tidal heating effect is considered. From the table, we find that the reflectivity parameter $|\mathcal{R}|^{2}$ of the ECO can be constrained within a fractional error of ${}10^{-6}$ via the observations of LISA. The measurement errors of the full waveform parameters is slightly decreased with the initial orbital eccentricity.
In most papers on EMRI data analysis and parameter estimation Babak:2009ua ; Wang:2012xh ; Ye:2024bku , the focus is primarily on the correlations among intrinsic parameters Katz:2021yft ; Zhang:2022rfr ; Zhang:2023vok . Therefore, in the following subsection, we only plot the correlations between the constraint reflectivity and the other intrinsic parameters. Fig. 4 shows the probability distribution for the mass of secondary body, mass of primary body, spin of primary body, initial orbital eccentricity, initial semi-latus rectum, initial orbital inclination and reflectivity $(M,m,a,e_{0},p_{0},x_{I},|\mathcal{R}|^{2})$, the results are plotted using the off-diagonal elements of the variance-covariance matrix. One can find that the bound on the reflectivity is closely associated with the parameters $(M,a,e_{0},p_{0},x_{I})$, and the correlation between the reflection $|\mathcal{R}|^{2}$ and secondary body mass $m$ is relative weaker.
Figure 4: The corner plot for the probability distribution for the mass of secondary body, mass of primary body, spin of primary body, initial orbital eccentricity, initial semi-latus rectum, initial orbital inclination and reflectivity $(M,m,a=0.9,e_{0}=0.4,p_{0}=12,x_{I},|\mathcal{R}|^{2})$, which is inferred from one-year observation of EMRI. Vertical lines show the $1\sigma$ interval for each source parameter. The contours correspond to $68\%$, $95\%$ and $99\%$ probability confidence intervals